OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14 i got b

OpenStudy (anonymous):

Its b

OpenStudy (anonymous):

On this part of the equation 14x - 6y = 4 The 4 is supposed to be 8

OpenStudy (anonymous):

And in this equation -14x + 28y = 1 The 1 is supposed to be 7

OpenStudy (anonymous):

-7*

OpenStudy (anonymous):

You need to factor everything out of the formula and see if you get the same answer, so: A) 4*(7x-3y=4) and -3*(2x-4y=1) B) 2*(7x-3y=2) and 1*(-14+28=1) C) -4*(7x-3y=4) and 14*(2x - 4y = 1) Then you can easily see that the answer is B, because there the factored form is not the same as the others.